# MAD-AD: Masked Diffusion for Unsupervised Brain Anomaly Detection

### v1

@@ 1 Introduction
The accurate detection and localization of brain anomalies in medical images, particularly in Magnetic Resonance Imaging (MRI) data, is paramount to diagnosing and understanding neurological injuries and pathologies. However, the complexity of brain structures and the scarcity of labeled abnormal data present significant challenges in developing robust and generalizable solutions. Traditionally, brain anomaly detection has been framed as a supervised learning task, which aims at identifying well-defined pathologies such as brain tumor [16, 18, 36], atrophy [29] or white matter hyper-intensities [20, 22], among many others. Nevertheless, casting anomaly detection as a supervised problem introduces an inherent bias towards the targeted lesions, limiting the scope of detectable pathologies. Moreover, collecting large amounts of annotated samples encompassing the entire spectrum of potential brain abnormalities is exceedingly expensive and is impractical for novel structures or rare abnormal patterns.
@@
Recent advances in diffusion models [17, 21, 37, 38] have opened new avenues in generative modeling. Such models [37] leverage a stochastic process to gradually corrupt data and learn to reverse this process, enabling them to model complex data distributions with remarkable precision. Their success in generating high-quality images and their ability to capture intricate patterns in the data have prompted researchers to explore their use for anomaly detection [15, 26, 40, 24, 6, 28]. While these methods have improved the accuracy of anomaly detection, their application to brain images introduces several challenges. Firstly, the forward diffusion process can cause a loss of distinctive features across brain regions, particularly when the number of steps is large. This loss, particularly when working with medical imaging data, may compromise the model’s ability to differentiate between normal and anomalous brain regions. Also, reducing the number of forward diffusion steps introduces the risk of an “identity shortcut” problem. In this problem, the model can easily recover the fine details of the input image, resulting in anomalous regions being preserved in the reconstruction. This is a significant concern in brain anomaly detection, where subtle but critical deviations such as tumors or lesions may be overlooked due to this shortcut behavior. Another issue arises from the indiscriminate application of forward and reverse diffusion across the entire brain image. This approach can hinder the model’s ability to effectively reconstruct normal brain patterns, especially when anomalies are localized. Discrepancies between the input and reconstructed images can lead to false positives in anomaly detection, especially in regions where the normal brain tissue is relatively homogeneous or where anomalies manifest as intricate or sparse patterns. Finally, when the input image contains random patterns or textures, such as in cases of neurological noise or motion artifacts, the diffusion process may struggle to recover the original structures, leading to false anomaly detections. This is particularly problematic in brain imaging, where the normal structures of the brain can have complex, interdependent patterns that are sensitive to minor perturbations.
@@ 2 Related works
AE-based methods.  Approaches in this category train a autoencoder on normal data to accurately reconstruct input images. At inference, the reconstruction error measured at each pixel is used to localize anomalies. Different networks have been explored for the reconstruction, including standard autoencoders (AE) [2, 4], variational autoencoders (VAEs) [4, 35, 42] and denoising autoencoders (DAEs) [19]. A common issue with these methods is their propensity to overfit the training data, leading to a poor generalization on unseen data. Furthermore, they are prone to blurry reconstructions, struggling to accurately distinguish subtle anomalies from normal variations, especially when relying solely on reconstruction error as a measure of abnormality.
@@
Diffusion-based method.  Diffusion models have gained significant attention in computer vision for their ability to generate high-fidelity images [13]. Recently, these models have also shown promise in various medical image analysis tasks including UAD [8, 6, 9, 10, 23, 27, 28, 39]. A prominent diffusion-based method for UAD in medical images, AnoDDPM [39] utilizes a partial diffusion strategy, adding noise to an image up to a specific timestep and then recovering the original image with a reverse diffusion process. This method has shown success in detecting anomalies in brain MRI and other domains. pDDPM [8] instead applies the diffusion process in a patch-wise manner, aiming to improve the understanding of local image context and achieve better anatomical coherence in the reconstruction. This method divides the image into overlapping patches and reconstructs each patch while considering its unperturbed surroundings. cDDPM [6, 9] leverages additional information from the input image, such as latent features obtained from an encoder, to further guide the denoising process. This conditioning aims to improve the reconstruction quality and the model’s ability to capture subtle anomalies. AutoDDPM [12] incorporates automatic masking, stitching, and resampling techniques within the DDPM framework to enhance its robustness and accuracy in anomaly detection. This approach also addresses the challenge of selecting an appropriate noise level for detecting lesions of various sizes.
@@
A recently proposed method, DISYRE [27, 28], uses a diffusion-like pipeline to train a model to restore images that have been corrupted with synthetic anomalies. Anomalies in a new image are detected based on the model’s ability to restore the image to a healthy state. A key limitation of this method is that the synthetic anomalies may not encompass all types of real-world anomalies, limiting its generalization ability. THOR [10] integrates implicit guidance into the DDPM’s denoising process using intermediate masks to preserve the integrity of healthy tissue details. It aims to ensure a faithful reconstruction of the original image in areas unaffected by pathology, minimizing false positives. However, since these intermediate masks are determined based on the perceptual differences between input images and their reconstruction at each step, the model may struggle to detect subtle or small anomalies, as they might be masked out due to their minimal differences with the input image. Additionally, reconstruction errors may occur due to the loss of details during the forward process, with normal regions not getting masked due to their high perceptual differences. IterMask2 [23] incorporates an iterative spatial mask refinement process and frequency masking to enhance UAD performance. This strategy minimizes information loss in normal areas by iteratively shrinking a spatial mask, starting from the whole brain towards the anomaly. Although the model performs well in detecting hypo- or hyper-intense areas, it can fail to localize structural anomalies such as atrophy or enlarged ventricles as their reconstruction is conditioned on structural information from high-frequency image components which can be recovered by the model.
@@ 3.1 Modeling the normal feature space
Random masking.  To incorporate random masking into the forward diffusion process, given the latent features of an input normal sample $\boldsymbol{z}_{0}\sim p\left(\boldsymbol{z}_{0}\right)$, we spatially divide $\boldsymbol{z}_{0}$ into non-overlapping patches defined by a random mask $M\in[0,1]^{H\times W}$. The forward Markov diffusion process to generate samples $\boldsymbol{z}_{t}$ gradually applies noise to the non-masked patches of sample $\boldsymbol{z}_{0}$ for $t$ time steps, where $t\in[1,T]$. Following [17], the forward noising process in the latent space with masking can be characterized as
@@
where $\boldsymbol{z}_{t}$ is the partially diffused image at step $t$, $\epsilon\sim\mathcal{N}(0,I)$ is the sampled Gaussian noise and $\beta_{t}$ is the noise schedule at step $t$, which controls the amount of noise added at each step. Using the reparameterization trick, $\boldsymbol{z}_{t}$ can be obtained implicitly as
@@
In this equation, $\mu_{\theta}(\boldsymbol{z}_{t},t)$ is a trainable function, which can be reparameterized as a predicted noise $\epsilon$ or a predicted clean image $\boldsymbol{z}_{0}$. Due to the incorporated random masking strategy, we prefer the latter one for simplicity. Therefore, $\mu_{\theta}(\boldsymbol{z}_{t},t)$ can be expressed as
@@
Mask prediction.  By parameterizing $f_{\theta}$ as a neural network, the model can be trained using a simple mean-square error loss between $\boldsymbol{z}_{0}$ and the predicted clean image. Moreover, in Eq. (3), we assumed that the mask $M$ is available in the reverse process. However, this assumption is unrealistic since the mask used in diffusing the image, which contains the location of anomalous regions, is not accessible at inference. Therefore, we include an additional head $f_{\theta,M}$ to the diffusion model that predicts the mask used in the forward diffusion. This can be achieved by applying a binary cross-entropy (BCE) loss between the predicted mask from this head and a randomly sampled mask used during partial diffusion in training. The final training objective of our model can be expressed as
$\min_{\theta}\ \ \mathbb{E}_{\boldsymbol{z}_{0}\sim q\left(\boldsymbol{z}_{0}\right),\boldsymbol{\epsilon},t,M}\left[\left\|\boldsymbol{z}_{0}-f_{\theta,\boldsymbol{z}_{0}}\left(\boldsymbol{z}_{t},t\right)\right\|_{2}^{2}\,+\,\lambda\!\cdot\!BCE\big{(}M,\,f_{\theta,M}(\boldsymbol{z}_{t},t)\big{)}\right],$ (5)
@@ 3.2 Recovering normal images
Nevertheless, one drawback of sampling with DDPM is that it requires many reverse sampling steps to obtain the normal version. Therefore, we instead opted for DDIM [38] which, by reducing the stochasticity of DDPM, makes the reverse process more deterministic and requires fewer sampling steps. Consequently, we modify the reverse process of DDIM for the MAD-AD model as
$\boldsymbol{z}^{\prime}_{t-1}=$ $\underbrace{B\Big{(}f_{\theta,M}(\boldsymbol{z}^{\prime}_{t})\Big{)}}_{\textrm{``predicted mask''}}\cdot\Big{(}\sqrt{\bar{\alpha}_{t-1}}\underbrace{f_{\theta,z_{0}}(z^{\prime}_{t})}_{\textrm{``predicted $\tilde{\boldsymbol{z}}^{\prime}_{0}$''}}+\underbrace{\sqrt{1-\bar{\alpha}_{t-1}}\tilde{\boldsymbol{\epsilon}}_{t}(\boldsymbol{z}^{\prime}_{t})}_{\textrm{``direction pointing to $\boldsymbol{z}^{\prime}_{t}$''}}+\ \sigma_{t}\boldsymbol{\epsilon}^{\prime}_{t}\Big{)}$ (6)
@@
Figure 2: Visual example of the reverse process. Both the predicted mask and $\boldsymbol{z}^{\prime}_{t}$ at multiple time steps are depicted to highlight the masked reverse sampling in MAD-AD.
@@ 3.3 Anomaly localization
Equation (6) enables a correcting trajectory from $\boldsymbol{z}^{\prime}_{T}$ to $\boldsymbol{z}^{\prime}_{0}$, resulting in generating high-quality normal variation of the anomalous image in fewer steps. To accurately localize anomalies, we used the discrepancy between the input image and its reconstructed normal counterpart. More concretely, using the “normal” latent embedding $\tilde{\boldsymbol{z}}^{\prime}_{0}$, we generated a reconstructed normal sample in the image-space as $\tilde{\boldsymbol{x}}^{\prime}_{0}=\mathcal{V}_{D,\phi}(\tilde{\boldsymbol{z}}^{\prime}_{0})$, where $\mathcal{V}_{D,\phi}$ is the pre-trained VAE decoder. The predicted anomaly map is then obtained as
$\boldsymbol{a}\,=\,G\ast\min(\|\tilde{\boldsymbol{x}}^{\prime}_{0}-\boldsymbol{x}^{\prime}_{0}\|,\gamma)/\gamma.$ (8)
@@ 4.2 Results
Method Pathology (Dice)$\uparrow$
@@
Method Modality
@@
Cycl.UNet [24]${}_{\tiny\textsl{MICCAI'23}}$ 65.0 42.6 49.5 37.0 48.5
@@
IterMask2 [23]${}_{\tiny\textsl{MICCAI'24}}$ 80.2 61.7 71.2 58.5 67.9
@@
Table 3: Effect of different sources for the anomaly score in MAD-AD (BRATS’24).
Anomaly source Modality
@@
Hyper- parameter Value Modality
@@ 5 Conclusion
This paper introduces a novel unsupervised anomaly detection method for brain MRI using a latent diffusion model with a random masking strategy. The approach leverages latent space, as brain anomalies in the latent space could be considered as noise and therefore be removed during the denoising process of diffusion models. Furthermore, by using a mask prediction module in the diffusion model, the model can selectively modify anomalous regions while preserving normal areas, enabling accurate identification of anomalous regions. Experiments on two datasets and two common brain UAD experimental settings demonstrate the superiority of our approach, validating its effectiveness in detecting and localizing brain anomalies without requiring labeled data, and showcasing is promising potential as alternative to existing methods.

### v2

@@ 1 Introduction
The accurate detection and localization of brain anomalies in medical images, particularly in Magnetic Resonance Imaging (MRI) data, is paramount to diagnosing and understanding neurological injuries and pathologies. However, the complexity of brain structures and the scarcity of labeled abnormal data present significant challenges in developing robust and generalizable solutions. Traditionally, brain anomaly detection has been framed as a supervised learning task, which aims at identifying well-defined pathologies such as brain tumor [15, 18, 36], atrophy [29] or white matter hyper-intensities [20, 22], among many others. Nevertheless, casting anomaly detection as a supervised problem introduces an inherent bias towards the targeted lesions, limiting the scope of detectable pathologies. Moreover, collecting large amounts of annotated samples encompassing the entire spectrum of potential brain abnormalities is expensive and impractical for novel structures or rare abnormal patterns.
@@
Recent advances in diffusion models [16, 21, 37, 38] have opened new avenues in generative modeling. Such models [37] leverage a stochastic process to gradually corrupt data and learn to reverse this process, enabling them to model complex data distributions with remarkable precision. Their success in generating high-quality images and their ability to capture intricate patterns in the data have prompted researchers to explore their use for anomaly detection [14, 26, 40, 23, 28]. While these methods have improved the accuracy of anomaly detection, their application to brain images introduces several challenges. Firstly, the forward diffusion process can cause a loss of distinctive features across brain regions, especially when the number of steps is large. This loss may compromise the model’s ability to differentiate between normal and anomalous brain regions. Also, reducing the number of forward diffusion steps introduces the risk of an “identity shortcut” problem. In this problem, the model can easily recover the fine details of the input image, resulting in anomalous regions being preserved in the reconstruction. This is a significant concern in brain anomaly detection, where subtle but critical deviations such as tumors or lesions may be overlooked due to this shortcut behavior. Another issue arises from the indiscriminate application of forward and reverse diffusion across the entire brain image. This approach can hinder the model’s ability to effectively reconstruct normal brain patterns.
@@ 2 Related works
AE-based methods.  Approaches in this category train an autoencoder on normal data to accurately reconstruct input images. At inference, the reconstruction error measured at each pixel is used to localize anomalies. Different networks have been explored for the reconstruction, including standard autoencoders (AE) [2, 4], variational autoencoders (VAEs) [4, 35, 42] and denoising autoencoders (DAEs) [19]. A common issue with these methods is their propensity to overfit the training data, leading to a poor generalization on unseen data. Furthermore, they are prone to blurry reconstructions, struggling to accurately distinguish subtle anomalies from normal variations, especially when relying solely on reconstruction error as a measure of abnormality.
@@
Diffusion-based method.  Diffusion models have gained significant attention in computer vision for their ability to generate high-fidelity images [12]. Recently, these models have also shown promise in various medical image analysis tasks including UAD [7, 17, 8, 9, 24, 27, 28, 39]. A prominent diffusion-based method for UAD in medical images, AnoDDPM [39] utilizes a partial diffusion strategy, adding noise to an image up to a specific timestep and then recovering the original image with a reverse diffusion process. This method has shown success in detecting anomalies in brain MRI and other domains. PDDPM [7] instead applies the diffusion process in a patch-wise manner, aiming to improve the understanding of local image context and achieve better anatomical coherence in the reconstruction. This method divides the image into overlapping patches and reconstructs each patch while considering its unperturbed surroundings. CDDPM [8] generates multiple reconstructions via the reverse diffusion process and pinpoints anomalies by examining the distribution of these reconstructions with the Mahalanobis distance, subsequently labeling outliers as anomalies. MDDPM[17] incorporates masking-based regularization, applied on both image patches and in the frequency domain, to enhance unsupervised anomaly detection. AutoDDPM [11] incorporates automatic masking, stitching, and resampling techniques within the DDPM framework to enhance its robustness and accuracy in anomaly detection. This approach also addresses the challenge of selecting an appropriate noise level for detecting lesions of various sizes. However, the diffusion-based UAD models mentioned above rely heavily on a forward diffusion process that inherently results in information loss. Consequently, these methods often fail to accurately reconstruct the original healthy brain structures, leading to false-positive detections where normal regions are incorrectly identified as anomalous. This issue is particularly prominent in brain anomaly detection tasks, as brain structures, especially cortical regions, vary uniquely across individuals, thereby increasing the difficulty of accurately recovering normal anatomical variations.
@@
A recently proposed method, DISYRE [27, 28], uses a diffusion-like pipeline to train a model to restore images that have been corrupted with synthetic anomalies. Anomalies in a new image are detected based on the model’s ability to restore the image to a healthy state. A key limitation of this method is that the synthetic anomalies may not encompass all types of real-world anomalies, limiting its generalization ability. THOR [9] integrates implicit guidance into the DDPM’s denoising process using intermediate masks to preserve the integrity of healthy tissue details. It aims to ensure a faithful reconstruction of the original image in areas unaffected by pathology, minimizing false positives. However, since these intermediate masks are determined based on the perceptual differences between input images and their reconstruction at each step, the model may struggle to detect subtle or small anomalies, as they might be masked out due to their minimal differences with the input image. Additionally, reconstruction errors may occur due to the loss of details during the forward process, with normal regions not getting masked due to their high perceptual differences. Inspired by diffusion-based models, IterMask2 [24] incorporates an iterative spatial mask refinement process and frequency masking to enhance UAD performance. This strategy minimizes information loss in normal areas by iteratively shrinking a spatial mask, starting from the whole brain towards the anomaly. Although the model performs well in detecting hypo- or hyper-intense areas, it can fail to localize structural anomalies such as atrophy or enlarged ventricles as their reconstruction is conditioned on structural information from high-frequency image components which can be recovered by the model.
@@ 3.1 Modeling the normal feature space
Random masking.  To incorporate random masking into the forward diffusion process, given the latent features of an input normal sample $\boldsymbol{z}_{0}\sim p\left(\boldsymbol{z}_{0}\right)$, we spatially divide $\boldsymbol{z}_{0}$ into non-overlapping patches defined by a random mask $M\in[0,1]^{H\times W}$. The forward Markov diffusion process to generate samples $\boldsymbol{z}_{t}$ gradually applies noise to the non-masked patches of sample $\boldsymbol{z}_{0}$ for $t$ time steps, where $t\in[1,T]$. Following [16], the forward noising process in the latent space with masking can be characterized as:
@@
where $\boldsymbol{z}_{t}$ is the partially diffused image at step $t$, $\epsilon\sim\mathcal{N}(0,I)$ is the sampled Gaussian noise and $\beta_{t}$ is the noise schedule at step $t$, which controls the amount of noise added at each step. Using the reparameterization trick, $\boldsymbol{z}_{t}$ can be obtained implicitly using the following equation:
@@
In this equation, $\mu_{\theta}(\boldsymbol{z}_{t},t)$ is a trainable function, which can be reparameterized as a predicted noise $\epsilon$ or a predicted clean image $\boldsymbol{z}_{0}$. Due to the incorporated random masking strategy, we prefer the latter one for simplicity. Therefore, $\mu_{\theta}(\boldsymbol{z}_{t},t)$ can be formally expressed as:
@@
Mask prediction.  By parameterizing $f_{\theta}$ as a neural network, the model can be trained using a simple mean-square error loss between $\boldsymbol{z}_{0}$ and the predicted clean image. Moreover, in Eq. (3), we assumed that the mask $M$ is available in the reverse process. However, this assumption is unrealistic since the mask used in diffusing the image, which contains the location of anomalous regions, is not accessible at inference. Therefore, we include an additional head $f_{\theta,M}$ to the diffusion model that predicts the mask used in the forward diffusion. This can be achieved by applying a binary cross-entropy ($\mathcal{L}_{BCE}$) loss between the predicted mask from this head and a randomly sampled mask used during partial diffusion in training. The final training objective of our model is defined by:
$\min_{\theta}\ \ \mathbb{E}_{\boldsymbol{z}_{0}\sim q\left(\boldsymbol{z}_{0}\right),\boldsymbol{\epsilon},t,M}\left[\left\|\boldsymbol{z}_{0}-f_{\theta,\boldsymbol{z}_{0}}\left(\boldsymbol{z}_{t},t\right)\right\|_{2}^{2}\,+\,\lambda\mathcal{L}_{BCE}\big{(}M,\,f_{\theta,M}(\boldsymbol{z}_{t},t)\big{)}\right],$ (5)
@@ 3.2 Recovering normal images
Nevertheless, one drawback of sampling with DDPM is that it requires many reverse sampling steps to obtain the normal version. Therefore, we instead opted for DDIM [38] which, by reducing the stochasticity of DDPM, makes the reverse process more deterministic and requires fewer sampling steps. Consequently, we modify the reverse process of DDIM for the MAD-AD model as:
$\boldsymbol{z}^{\prime}_{t-1}=$ $\underbrace{B\Big{(}f_{\theta,M}(\boldsymbol{z}^{\prime}_{t})\Big{)}}_{\textrm{``predicted mask''}}\Big{(}\sqrt{\bar{\alpha}_{t-1}}\underbrace{f_{\theta,z_{0}}(z^{\prime}_{t})}_{\textrm{``predicted $\tilde{\boldsymbol{z}}^{\prime}_{0}$''}}+\underbrace{\sqrt{1-\bar{\alpha}_{t-1}}\tilde{\boldsymbol{\epsilon}}_{t}(\boldsymbol{z}^{\prime}_{t})}_{\textrm{``direction pointing to $\boldsymbol{z}^{\prime}_{t}$''}}+\ \sigma_{t}\boldsymbol{\epsilon}^{\prime}_{t}\Big{)}$ (6)
@@
Figure 2: Visual example of the reverse process. Both the predicted mask and the decoded latent representation of the intermediate reverse step ($\boldsymbol{z}^{\prime}_{t}$) at multiple time steps are depicted to highlight the masked reverse sampling in MAD-AD.
@@ 3.3 Anomaly localization
Equation 6 enables a correcting trajectory from $\boldsymbol{z}^{\prime}_{T}$ to $\boldsymbol{z}^{\prime}_{0}$, resulting in generating high-quality normal variation of the anomalous image in fewer steps. To accurately localize anomalies, we used the discrepancy between the input image and its reconstructed normal counterpart. More concretely, using the “normal” latent embedding $\tilde{\boldsymbol{z}}^{\prime}_{0}$, we generated a reconstructed normal sample in the image-space as $\tilde{\boldsymbol{x}}^{\prime}_{0}=\mathcal{V}_{D,\phi}(\tilde{\boldsymbol{z}}^{\prime}_{0})$, where $\mathcal{V}_{D,\phi}$ is the pre-trained VAE decoder. The predicted anomaly map is then given by:
$\boldsymbol{a}\,=\,G\ast\min(\|\tilde{\boldsymbol{x}}^{\prime}_{0}-\boldsymbol{x}^{\prime}_{0}\|,\gamma)/\gamma,$ (8)
@@ 4.2 Results
Method Pathology (Global Max Dice)$\uparrow$
@@
Method Modality (Max Dice)$\uparrow$
@@
Cycl.UNet [${}_{\tiny\textsl{MICCAI'23}}$ 65.0 42.6 49.5 37.0 48.5
@@
IterMask2 [${}_{\tiny\textsl{MICCAI'24}}$ 80.2 61.7 71.2 58.5 67.9
@@
Table 3: Effect of different sources for the anomaly score in MAD-AD (BRATS’21).
Anomaly source Modality (Max Dice)$\uparrow$
@@
Hyper- parameter Value Modality (Max Dice)$\uparrow$
@@ 5 Conclusion
This paper introduces a novel unsupervised anomaly detection method for brain MRI using a latent diffusion model with a random masking strategy. The approach leverages latent space, as brain anomalies in the latent space could be considered as noise and therefore be removed during the denoising process of diffusion models. Furthermore, by using a mask prediction module in the diffusion model, the model can selectively modify anomalous regions while preserving normal areas, enabling accurate identification of anomalous regions. Experiments on two datasets and two common brain UAD experimental settings demonstrate the superiority of our approach, validating its effectiveness in detecting and localizing brain anomalies without requiring labeled data, and showcasing its promising potential as an alternative to existing methods.